OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 3x + 2y = 6 2x + 3y = -1 Answer -3 4 -4 3

OpenStudy (anonymous):

I am going to solve by substitution but elimination will work too 3x + 2y = 6 (1) 2y=6-3x y=3-(3/2)x 2x + 3y = -1 (2) Sub y from (1) into (2) 2x+3(3-(3/2)x)=-1 2x+9-(9/2)x=-1 (-5/2)x=-10 -5x=-20 x=4